OpenStudy (goldrush18):

Solve the simultaneous equations x^2+xy=6 x-3y+1=0

OpenStudy (anonymous):

substitute y from the second equation in the first equation. the first equation becomes a quadratic in x then you'll be easily able to solve it

OpenStudy (goldrush18):

im trying but still don't know if i'm right

OpenStudy (anonymous):

These are the answers after following @vamgadu's recommended procedure:\[\left\{\left\{x=-\frac{9}{4},y=-\frac{5}{12}\right\},\{x=2,y=1\}\right\} \]

OpenStudy (goldrush18):

ok thanks but i'm stuck at the part where i substitute y= -1-x/-3 into the first equation